# Exergetic Port-Hamiltonian Systems Modeling Language

### v1

@@ 10 Discussion
While it still seems reasonable to package a motor model as a single bond graph, integrating the model as part of a more complex system is not supported by the bond graph notation. Hence, bond graph users have to reproduce models as parts of more complex models. EPHS improves upon this, making it nearly trivial to encapsulate, reuse, and swap out submodels. Another main feature of EPHS is their built-in thermodynamic consistency. Restricting the space of models in this way not only provides important guarantees, it also makes models compatible with each other, as a close link between their mathematical structure and physical interpretation is established. This is reflected by the correspondence of junctions and energy domains, which in our opinion contributes substantially to the improved readability of the EPHS syntax. The reason why EPHS users can easily integrate existing models without first needing to adapt them to fit into the larger system is that EPHS simply compose by sharing energy domains. In this way, we hope to improve upon port-Hamiltonian systems, which lack an intuitive graphical syntax and require that users define the ports of a system in a somewhat arbitrary manner.
@@
Markus Lohmayer: Conceptualization, Investigation, Writing – Original Draft, Writing – Review & Editing, Visualization; Owen Lynch: Investigation, Writing – Original Draft; Sigrid Leyendecker: Supervision
Acknowledgements
We thank David Spivak for help when working out the concept of namespaces.

### v2

@@ 10 Discussion
10.1 Comparison with bond graphs
Bond graphs provide a graphical notation for representing physical systems in terms of subsystems that exchange energy. However, the notation does not naturally support the hierarchical nesting of systems. Relying exclusively on primitive subsystems limits the applicability of bond graphs to complex systems, where a hierarchical breakdown into simpler parts would facilitate both model construction and reuse.
Although the notation is not inherently composable, one can introduce conventions that allow for subsystems defined by open bond graphs. For example, The convention in [50] considers bond graphs that share a specific element in common as a modules. These modules can be ‘glued’ together in a parent-level bond graph via an additional 00-junction. In this approach, the common element in each individual bond graph is replaced by an external port; the element is instantiated once in a parent-level bond graph, where it is connected to a 00-junction. The modules are represented by additional elements, and their ports are also connected to the 00-junction. The direction of these bonds must match the direction of the corresponding bond (or port) at the module level.
While port-Hamiltonian systems can be seen as a geometric formalization of the dynamical systems derived from generalized bond graphs, the algebraic formalization of the bond graph notation itself is not straightforward (see, e.g., [51, 32]). Some inherent complexity in the notation is evident also in practice, where the direction in which exchange of power is counted as positive, and sometimes also the role of a variable as flow versus effort, must be decided on a bond-by-bond basis. These decisions must either be left to the user or managed through heuristic algorithms or conventions. This can impede seamless hierarchical decomposition and reuse of subsystems.
Interconnection patterns provide a composable, graphical syntax for the EPHS modeling language. Compared to bond graphs, interconnection patterns have an outer box, and they exclusively rely on 00-junctions. Based on this simplification, the mentioned issues are resolved:
  • 1.
    Interconnection patterns have a clear mathematical structure, which supports their hierarchical (de-)composition.
  • 2.
    Interconnection patterns are relatively easy to interpret, owing to their simple structure, the direct correspondence between junctions and energy domains, and the clear role of a power variables as either flow or effort.
  • 3.
    The direction in which exchange of power is counted positive is predetermined in agreement with the convention used in thermodynamics, i.e. energy supplied to a system has a positive sign.
In contrast to bond graphs, EPHS models are guaranteed to be thermodynamically consistent.
Bond graphs may incorporate elements from block diagrams. In addition to bonds, which represent energy exchange, there may be edges that transmit signals according to their orientation. Elements may have signal ports to receive signals, which modulate their behavior, and there may be ‘blocks’ that perform arbitrary signal processing. In contrast, the formalism presented here does not cover arbitrary signal processing; system behavior can be modulated only by state variables of other systems. To express this, every port has a state variable, which is shared with all other ports connected to the same junction. We distinguish between state ports, which share only state, and power ports, which additionally model energy exchange.
10.2 Comparison with port-Hamiltonian systems
Port-Hamiltonian Systems (PHS) provide a mathematical framework for open dynamical systems with guaranteed passivity. While bond graphs are frequently used to depict port-Hamiltonian models, establishing a rigorous link between bond graph syntax and port-Hamiltonian semantics remains challenging. On the one hand, bond graphs offer a network representation of individual systems, but the representation is not inherently composable. On the other hand, PHS provide a composable representation of individual systems, albeit without an explicit internal network structure. In particular, the PHS framework provides no built-in support for referencing subsystems and their ports. Previous work has focused on algorithms for converting certain classes of bond graphs to PHS (see, e.g., [52]).
A well-known issue concerns the relationship between port-Hamiltonian structure, which implies passivity, and its physical interpretation. Owing to its roots in mechanics, the ‘Hamiltonian’ storage function is typically interpreted as the system’s energy. However, the differential of the ‘Hamiltonian’ generally also drives dissipative dynamics, resulting in an apparent loss of ‘energy’. This raises questions, especially outside the realm of electromechanical systems, where thermodynamic considerations are often overlooked. In response, some members of the port-Hamiltonian community have proposed alternative port-based frameworks that explicitly encode thermodynamic structure (see, e.g., [53, 54, 55]).
EPHS provide a mathematical framework for open dynamical systems with guaranteed thermodynamic consistency. In contrast to PHS, this framework is directly based on a composable, graphical syntax used to specify how more complex systems are formed by interconnecting simpler systems. An explicit understanding of the translation between graphical syntax and relational semantics guarantees that syntax and semantics compose in compatible ways. This enables a modular and hierarchical modeling approach, which inherently supports referencing nested subsystems and their ports. In a sense, this approach unites the strengths of both bond graphs and PHS.
The dynamical systems defined by EPHS models essentially possess port-Hamiltonian structure. Beyond passivity, additional structural properties imposed on the primitive systems guarantee that all EPHS share a uniform and consistent physical interpretation, aligned with nonequilibrium thermodynamics. Since EPHS compose simply by sharing energy domains, users typically do not need to explicitly define external ports. As a result, users can typically integrate existing models into more complex systems without any prior adaptation.
10.3 Comparison with metriplectic and GENERIC systems
Metriplectic systems and the GENERIC provide a framework for (usually isolated) dynamical systems with guaranteed thermodynamic consistency. The dynamical systems defined by (isolated) EPHS models are basically metriplectic systems or instances of the GENERIC, since the built-in thermodynamic consistency follows from essentially the same ideas. Regarding the reversible part, EPHS appear to be more general, since they use Dirac structures, rather than the less general Poisson structures used in the definition of metriplectic systems or the GENERIC. This allows EPHS to easily model kinematic constraints, which is important e.g. for multibody systems (see [56]). However, it should be no problem to generalize the metriplectic or GENERIC framework accordingly. Regarding the irreversible part, EPHS appears to be less general, since it remains unclear if the single-generator formulation based on exergy is useful beyond local thermodynamic equilibrium.
In contrast to the metriplectic or GENERIC framework, EPHS enable a modular and hierarchical approach to thermodynamic modeling, based on a graphical syntax for building more complex models from simpler parts.
10.4 Future work
Future work shall address the details omitted in this work. A more technical paper will define interconnection patterns and their functorial semantics using a categorical framework based on directories and the theory of operads/multicategories.
As presented here, the definition of (composite) EPHS exhibits a recursive structure. It would be interesting to study the ‘closure’ of all EPHS models by deriving representations of EPHS as structured systems of differential-algebraic equations. The analysis could benefit further developments and further clarify the relationship with the metriplectic or GENERIC framework as well as port-Hamiltonian systems.
As an important groundwork for further practical developments, we have implemented the framework within the Julia programming language, see [57]. The reader is invited to review its application in several examples. Based on this, future work can explore topics such as advanced applications (see, e.g., [56, 11]), structure-preserving numerical methods, semi-automated control design, and the integration with scientific machine learning methods.
@@ 10.4 Future work
Markus Lohmayer: Conceptualization, Investigation, Writing – Original Draft, Writing – Review & Editing, Visualization; Owen Lynch: Investigation, Writing – Original Draft; Sigrid Leyendecker: Supervision, Review & Editing, Funding
Acknowledgments
We thank David Spivak for help when working out the concept of namespaces. The second author acknowledges support from the DARPA ASKEM program under grant HR00112220038.